# Graph Transformers: A Survey

### v1

@@ VI-B Generalization and Robustness
    Applying domain adaptation or transfer learning techniques to facilitate learning from multiple source domains and transfer the knowledge from source domains to target domains.
@@
    Exploring meta-learning or few-shot learning techniques to enable learning from limited data and rapid adaptation to new tasks.
@@
VI-C Interpretability and Explainability
Graph transformers commonly regarded as black box models present significant challenges in terms of interpretability and explainability. The lack of sufficient justification and evidence for their decisions can undermine their credibility and transparency. To address this issue, several approaches can be considered:
@@ VI-C Interpretability and Explainability
    Applying explainable artificial intelligence (XAI) techniques, such as saliency maps, influence functions, and counterfactual explanations, to analyze and understand the behaviour and logic of graph transformer models.
@@
    Exploring natural language generation techniques, such as template-based generation, neural text generation, and question-answering generation, to produce natural language explanations for outputs or actions of graph transformer models.
@@
VI-D Learning on Dynamic Graphs
Graphs, which are frequently characterized by their dynamic and intricate nature, possess the ability to transform over time as a result of the addition or removal of nodes and edges, as well as the modification of node and edge attributes [206]. Moreover, these graphs may exhibit diverse types and modalities of nodes and edges. In order to empower graph transformers to effectively manage such dynamic graphs, it is advisable to explore the following strategies:
@@ VI-D Learning on Dynamic Graphs
    Developing temporal and causal attention mechanisms, such as recurrent, temporal, and causal attention, to capture the temporal and causal evolution of graphs.
@@
    Applying continual learning techniques on dynamic graphs to void forgetting previous knowledge and retraining.
@@
VI-E Data Quality and Diversity
Graph transformers require a significant quantity of diverse and high-quality data to achieve effective learning. Nevertheless, in the real world, data is frequently limited, noisy, incomplete, imbalanced, and biased. This has a detrimental effect on the performance and fairness of graph transformer models. To mitigate these challenges related to data quality and diversity, several potential strategies can be considered:
  1. 1.
    Developing noise-tolerant attention mechanisms, such as denoising attention, error correction attention, and self-supervised attention, to filter out data noise and errors.
  2. 2.
    Applying data augmentation or generation techniques, such as graph augmentation [207], graph completion, and graph generation, to enhance the quantity and variety of data.
  3. 3.
    Exploring data imbalance and bias mitigation techniques, such as data re-sampling, data re-weighting, and data debiasing, to improve data fairness and equity.
  4. 4.
    Evaluating data quality and diversity using data quality assessment or data diversity analysis methods, such as data quality metrics, data diversity measures, and data quality visualization.
VII Conclusion
Graph transformers are a novel and powerful class of neural network models, which can effectively encode and process graph-structured data. This survey provides a comprehensive overview of graph transformers in terms of design perspectives, taxonomy, applications, and open issues. We first discuss how graph transformers incorporate graph inductive bias, including node positional encodings, edge structural encodings, message-passing bias and attention bias, to encode the structural information of graphs. Then, we introduce the design of graph attention mechanisms, including global and local attention mechanisms. Afterwards, a taxonomy of graph transformers is presented. This survey also includes a design guide for effective graph transformers, including the best practices and recommendations for selecting appropriate components and hyperparameters. Moreover, the application scenarios of graph transformers are reviewed based on various graph-related tasks (e.g., node-level, edge-level, and graph-level tasks), as well as tasks in other domains. Lastly, current challenges and future directions of graph transformers are identified. This survey aims to serve as a valuable reference for researchers and practitioners interested in graph transformers and their applications.
Acknowledgment
The authors would like to thank Estrid He (RMIT University) for her helpful comments.

### v2

@@ VII-B Generalization and Robustness
    Applying domain adaptation or transfer learning techniques to learn from multiple source domains and transfer knowledge to target domains.
@@
    Exploring meta-learning or few-shot learning techniques for learning from limited data and rapid adaptation to new tasks.
@@
VII-C Interpretability and Explainability
Graph transformers, often seen as black box models, face challenges in interpretability and explainability, undermining their credibility and transparency. To address this, several approaches can be considered:
@@ VII-C Interpretability and Explainability
    Applying explainable artificial intelligence (XAI) techniques, such as saliency maps, influence functions, and counterfactual explanations, to analyze and understand the behavior and logic of graph transformer models.
@@
    Exploring natural language generation techniques, such as template-based generation, neural text generation, and question-answering generation [207], to produce natural language explanations for outputs or actions of graph transformer models.
@@
VII-D Learning on Dynamic Graphs
Graphs are dynamic and complex, often changing over time with the addition or removal of nodes and edges, and the modification of their attributes [208, 209]. They may also have diverse types and modalities of nodes and edges. To enable graph transformers to effectively handle such dynamic graphs, consider the following strategies:
@@ VII-D Learning on Dynamic Graphs
    Developing temporal and causal attention mechanisms, such as recurrent, temporal, and causal attention, to capture the temporal and causal evolution of graphs [210].
@@
    Applying continual learning techniques on dynamic graphs to avoid forgetting previous knowledge and retraining.
@@
VIII Conclusion
Graph transformers are a novel and powerful class of neural network models, which can effectively encode and process graph-structured data. This survey provides a comprehensive overview of graph transformers in terms of design perspectives, taxonomy, applications, and open issues. We first discuss how graph transformers incorporate graph inductive bias, including node positional encoding, edge structural encoding, message-passing bias and attention bias, to encode the structural information of graphs. Then, we introduce the design of graph attention mechanisms, including global and local attention mechanisms. Afterwards, a taxonomy of graph transformers is presented. This survey also includes a design guide for effective graph transformers, including the best practices and recommendations for selecting appropriate components and hyperparameters. Moreover, the application scenarios of graph transformers are reviewed based on various graph-related tasks (e.g., node-level, edge-level, and graph-level tasks), as well as tasks in other domains. Lastly, current challenges and future directions of graph transformers are identified. This survey aims to serve as a valuable reference for researchers and practitioners interested in graph transformers and their applications.